OpenStudy (maram):

Use factoring, the quadratic formula, or identities to solve the equation. Find the solution in the interval [0, 2π). 4 sin^2(x) + 5 sin (x) = 9

OpenStudy (anonymous):

\[ 4 u^2 - 5 u - 9=(1+u)(-9+4u)\\ 4 \sin^2(x) - 5 \sin(x) -9=( 1+ \sin(x))(-9 + 4 \sin(x))\\ ( 1+ \sin(x))(-9 + 4 \sin(x))=0\\ \] You can finish it now.

OpenStudy (maram):

thank you